OpenStudy (anonymous):

Solve for y: 2/3y + 8 = 6 Answer y = 3 y = -5 y = 5 y = -3 i got y = 3

jimthompson5910 (jim_thompson5910):

2/3y + 8 = 6 2/3y + 8 - 8= 6 - 8 2/3y = -2 (2/3y)*3 = -2*3 2y = -6 2y/2 = -6/2 y = -3 So the solution is y = -3